OpenStudy (anonymous):

what is the square root of 50 plus the square root of 50

OpenStudy (loser66):

=2squaroot of 50

OpenStudy (akashdeepdeb):

2 square root of 50 :D

OpenStudy (austinl):

\(\large{\sqrt{50}+\sqrt{50}=2\sqrt{50}}=10\sqrt{2}\)